what are the next three fractions in the pattern???
1/12, 1/6 1/4, 1/3, 5/12....

Answer by jim_thompson5910(35256)

hint: rewrite each fraction with the LCD 12

1/12 = 1/12 (no change needed)

1/6 = 2/12

1/4 = 3/12

1/3 = 4/12

5/12 = 5/12 (no change needed)